OpenStudy (anonymous):

A ramp has an angle of elevation of 19 degrees, and a length of 35 ft. What is the height in inches of the ramp at its high end?

OpenStudy (jdoe0001):

are you doing trig?

OpenStudy (anonymous):

I suppose so, but I'm learning it in geometry.

OpenStudy (jdoe0001):

well, you have an angle, and a SIDE, so, that'd imply the trig function SINE, or \( sin(\theta)=\cfrac{opp\ side}{hypothenuse} \implies opp\ side=sin(\theta)\times hypothenuse \)

OpenStudy (anonymous):

Thanks so much!

OpenStudy (jdoe0001):

yw